# MM-Prompt: Cross-Modal Prompt Tuning for Continual Visual Question Answering

### v1

@@ 3.4 Training Objective
where $\mathcal{L}_{\text{CE}}$ is the cross entropy loss that optimizes task performance. $\mathcal{L}_{\text{qk-align}}=\sum\nolimits_{i\in\mathcal{I}^{M}}^{M\in\{\text{Q},\text{V}\}}(1-\text{cos}(\mathbf{q}^{M},\mathbf{k}^{M}_{i}))$ maintains structural quality of our representations through query-key alignment [10, 11]. $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ ensure consistency between recovered representations during both intra-modal recovery and inter-modal recovery phases. This multi-component loss jointly optimizes for predictive accuracy, cross-modal alignment, and recovery quality, creating a balanced optimization objective that effectively enhances the quality and utility of cross-modal prompts.
As illustrated in Fig. 1, our MM-Prompt model differs fundamentally from existing CVQA approaches [6, 17, 4, 9, 10, 11, 13, 12] by introducing explicit cross-modal interactions at multiple stages. The overall workflow, shown in Fig. 2, consists of two complementary mechanisms that work together to maintain balanced modality representation. First, input features from vision and question modalities ($\mathbf{F}^{\text{V}}$ and $\mathbf{F}^{\text{Q}}$) undergo Cross-Modal Prompts Query. In this stage, each modality’s features attend to the opposite modality before generating query vectors. These enriched queries are then used to select relevant prompts. Then, a mask will be applied on the weighted sum of these selected prompts to create explicit pathways for cross-modal prompts recovery. The masked prompts then first process through intra-modal recovery to establish basic modality-specific patterns while introducing light cross-modal influence, followed by cross-modal that further integrates information across modalities through attention mechanisms and selective enhancement. The result is a set of recovered prompts that maintain balanced representations from both modalities.
@@ 4.1 Set Up
Dataset. Following [4], we conduct experiments on two widely-used datasets: VQA v2 [19] and NExT-QA [39]. VQA v2 contains 1.1 million pairs of real-world images and human-written questions, while NExT-QA includes 52K annotated video-based question-answer pairs. For VQA v2, we structure the continual learning experiments into 8 sequential tasks for DI and CI, while 10 sequential tasks for QI. In QI, each task introduces 1 new question type. For CI, each task contains 10 unique object classes. Under DI, each task contains 5 subtasks and each subtask focuses on 2 different question types about the same 10 object classes. For NExT-QA, we structure the experiments into 7 sequential tasks. For QI, each task introduces 1 new question type. For CI, 4 tasks contain 11 object classes each, and 3 tasks contain 12 object classes each. For DI, each task contains unique question types with 5 subtasks, where each subtask addresses the same 16 object classes.
@@
Implementation Details. We construct MM-Prompt as illustrated in Fig. 2. The pretrained transformer backbone consists of 12 stacked blocks for both encoder and decoder, with each attention layer containing 12 attention heads. All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory. Following DualPrompt [10], we further divide our question and visual prompts into General (G) and Expert (E) types that target different transformer layers. We also include a weight decay [40] on the $\mathcal{L}_{\text{inter}}$ to prevent degeneration of this alignment mechanism.
@@ 4.2 Experimental Results
In this section, we investigate and evaluate MM-Prompt against 9 outstanding and state-of-the-art methods, including 6 general CL methods [10, 12, 11, 17, 13] and 3 CVQA methods [6, 4, 9]. For fair comparison, all models contain at least two types of prompts(Q and V) except L2P* and use approximately the same number of prompts with identical injection layers if possible.
Based on Table 1, we draw several important conclusions: (1) Methods adopting cross-modal prompt isolation like Dual Prompt [10] and L2P [11] consistently show higher forgetting rates and lower accuracy across settings, confirming that isolated prompt selection and injection lead to degradation in performance due to imbalanced representations. (2) MM-Prompt consistently outperforms all the other methods across all settings. On VQA v2, it achieves 36.223% in DI (vs. MaPLe’s 35.187%), 39.240% in CI (vs. MaPLe’s 37.450%), and 16.757% in QI (vs. VQACL’s 15.525%). Notably, it reduces forgetting to less than half of the next best method in DI. On NExT-QA, MM-Prompt similarly outperforms in accuracy and maintains the lowest forgetting rates. While the second-best method varies across different incremental scenarios, MM-Prompt’s consistent superiority highlights its robustness across diverse settings. (3) We observe striking dataset-dependent difficulty patterns: VQA v2 (static images) shows poorest performance on QI, while NExT-QA (videos) struggles most with CI, suggesting that temporal dynamics fundamentally alter how modality conflicts during continual learning. (4) Among comparison methods, MaPLe [17] performs relatively well, which we attribute to its gradient-based visual-linguistic interactions. (5) Specialized CVQA methods like Triplet [6] and VQACL [4] underperform despite their success in classification-based settings and with memory buffers, highlighting MM-Prompt’s efficiency for maintaining balanced representations across modalities. These results collectively validate our approach’s effectiveness in maintaining balanced cross-modal representations across diverse incremental learning scenarios.
@@ 4.3 Modality Integration Analysis
To demonstrate that isolated prompts amplify modality imbalance, we train the models as normal and evaluate them on joint, vision-only and question-only inputs, then compute Modality Merge Effectiveness and Modality Difference, details for these settings are in Appendix A. As shown in Fig. 4, existing prompt-based methods [10, 11, 4] that adopt cross modal prompts isolation show low merge effectiveness and high modality gaps during the continual learning process. In contrast, our method maintains consistently high merge effectiveness and minimal modality disparity across new tasks, while achieving superior overall accuracy. These improvements stem from our two-component design: cross-modal prompt query prevents selection bias by enriching queries with complementary information, while cross-modal prompt recovery creates explicit pathways for modality interaction and fulfill them with cross-modal information. Together, these mechanisms counteract modality imbalance at critical points where isolation typically occurs in existing methods.
@@
Figure 5: Bottom up attention visualization[41] during inference on Dual Prompt [10] and MM-Prompt.
(a) Dual Prompt[10] ($A:\text{33.601},F_{\text{inter}}:2.660$)
(b) MM-Prompt ($A:\text{36.223},F_{\text{inter}}:\text{0.447}$)
@@
Figure 6: Visualization of injected prompts and input feature using t-SNE [42].
@@
Figs. 5 and 6 provide further qualitative evidence. We visualize the attention from the model on the bounding boxes during the inference using [41]. The attention maps (Fig. 5) show MM-Prompt precisely focuses on relevant image regions (correctly identifying “broccoli” and “airport”), while existing methods such as Dual Prompt produce incorrect answers with diffuse attention patterns, indicating possible occurrence of modality bias [18]. The t-SNE visualization [42] of injected prompts (Fig. 6) confirms the cross-modal merging effectiveness of MM-Prompt. As shown in Fig. 6(b), MM-Prompt’s prompts form more tightly assembled clusters compared to Dual Prompt in Fig. 6(a), where prompts remain clearly separated by modality. This observation demonstrates how our cross-modal prompt query and recover mechanisms help prompts encode integrated information across modalities rather than remaining isolated within modality-specific subspaces. Additionally, examining the input features from the question “What green vegetable is on the plate?” in Fig. 5 reveals another key difference: in MM-Prompt, input features position at the intersection between Q and V prompt clusters, while prompts in Dual Prompt remain distinctly separated by modality. This demonstrates that MM-Prompt’s injected prompts contains shared representational properties that facilitate effective cross-modal understanding and prevent modality-isolated prompts throughout sequential learning.
@@ 4.4 Ablation Study
CQ CR $A$ ($\uparrow$) $F_{\text{inter}}(\downarrow)$
@@
Performance with Memory Buffer For NExT-QA dataset, which contains complex video question-answering tasks with longer sequences and fewer training data, we also evaluate models with a memory buffer of 500 samples (Table 3). As a result, MM-Prompt substantially outperforms all the other methods, given that all methods benefit from the same memory capacity. These results confirm that MM-Prompt’s approach to create explicit pathways for modality interaction remain effective when supplemented with traditional continual learning techniques [26, 27].
Impact of Task Order. Figs. 8(a) and 8(b) demonstrate MM-Prompt’s stability across different learning scenarios. MM-Prompt maintains more consistent performance across different task orders, while Dual Prompt[10] exhibits more varied performance. By maintaining balanced representations through explicit cross-modal pathways in injected prompts, MM-Prompt reduces the cascading effects of task order biases that typically compound in existing prompt methods. This order-invariant behavior strongly evidences that MM-Prompt successfully maintains balanced representations throughout the continual learning process.
Computational Efficiency. Fig. 7 shows that the inference efficiency of MM-Prompt remains competitive with existing methods. Processing 100 samples takes 0.179 seconds, which is more efficient than MAPLE [17] (0.211s) and TRIPLET [6] (0.226s), while only slightly slower than Dual Prompt [10] (0.158s). This modest computational overhead delivers substantial performance improvements and significant forgetting reduction compared to the Dual Prompt [10]. This demonstrates that fostering the cross-modal understanding can be achieved with reasonable computational costs, making MM-Prompt practical for real-world continual learning applications.
Methods NExT-QA
Mem $A$ ($\uparrow$) $F_{\text{inter}}$ ($\downarrow$)
Triplet [6] 500 27.996 3.282
Dual Prompt [10] 500 27.641 3.765
CODA [12] 500 30.632 2.471
Maple[17] 500 30.049 3.112
VQACL [4] 500 30.753 2.692
CluMo [9] 500 29.776 3.248
MM-Prompt 500 32.394 2.259
MM-Prompt 200 29.945 3.112
MM-Prompt 1000 34.268 2.427
Figure 7: Time spent processing 100 samples.
(a) Dual Prompt [10]
(b) MM-Prompt
@@ A.1 Dataset Partitioning
For VQA v2 [19], we organize the continual learning experiments into 8 sequential tasks for both Class Increment (CI) and Dual Increment (DI) settings. In the CI setting, each task introduces 10 unique object classes. In the DI setting, each task consists of 5 subtasks, each focusing on 2 distinct question types about the same 10 object classes. For the Question Increment (QI) setting, we define 10 sequential tasks, each introducing 1 new question type.
For NExT-QA [39], we structure the experiments into 7 sequential tasks due to the dataset’s scale. In the QI setting, each task introduces 1 new question type. In the CI setting, 4 tasks contain 11 object classes each, while the remaining 3 tasks contain 12 object classes each. In the DI setting, each task consists of 5 subtasks with unique question types, each subtask cover 16 different object classes.
@@ A.2 Partitioning Details and Group Definitions
For NExT-QA [39], we structure the experiments into 7 sequential tasks following the approach of [4]. In the DI setting, we implement the hierarchical structure defined in [4], with question increments ordered according to Table 8. For the QI setting, we utilize the question type ordering specified in Table 8 while maintaining consistent visual classes across all tasks. In the CI setting, we employ the object class groups outlined in Table 5 while preserving all question types throughout the tasks.
@@ B.1 Modality Merge Effectiveness and Modality Difference
Modality Merge $\text{Modality Merge Effectiveness}=\text{Acc}^{\text{joint}}-\max(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}}).$ (12)
@@
Modality $\text{Modality Difference}=\max(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}})-\min(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}}).$ (13)
@@
Question Types Recognition Location Judge Commonsense Count Action Color Type Subcategory Causal
@@ B.2 Insertion Layers and Number of Prompts
Number of Prompts $A$ ($\uparrow$) $F_{\text{inter}}$($\downarrow$)
@@
Following [10], we insert General prompts at lower layers {1,2} and Expert prompts at intermediate layers {3,4,5}. This configuration follows the hierarchical processing pattern of the transformer, where lower layers handle more specific, fine-grained patterns, while intermediate layers process more abstract, transferable representations.

### v2

@@ 3.4 Training Objective
where $\mathcal{L}_{\text{CE}}$ is the cross entropy loss that optimizes task performance. $\mathcal{L}_{\text{qk-align}}=\sum\nolimits_{i\in\mathcal{I}^{M}}^{M\in\{\text{Q},\text{V}\}}(1-\text{cos}(\mathbf{q}^{M},\mathbf{k}^{M}_{i}))$ maintains structural quality of our representations through query-key alignment (Wang et al., 2022a, b). $\mathcal{L}_{\text{inter}}$ and $\mathcal{L}_{\text{intra}}$ ensure consistency between recovered representations during both intra-modal recovery and inter-modal recovery phases. This multi-component loss jointly optimizes for predictive accuracy, cross-modal alignment, and recovery quality, creating a balanced optimization objective that effectively enhances the quality and utility of cross-modal prompts.
As illustrated in Fig. 1, our MM-Prompt model differs fundamentally from existing CVQA approaches Qian et al. (2023); Khattak et al. (2023); Zhang et al. (2023); Cai and Rostami (2024); Wang et al. (2022a, b); Menabue et al. (2024); Smith et al. (2023) by introducing explicit cross-modal interactions at multiple stages. The overall workflow, shown in Fig. 2, consists of two complementary mechanisms that work together to maintain balanced modality representation. First, input features from vision and question modalities ($\mathbf{F}^{\text{V}}$ and $\mathbf{F}^{\text{Q}}$) undergo Cross-Modal Prompts Query. In this stage, each modality’s features attend to the opposite modality before generating query vectors. These enriched queries are then used to select relevant prompts. Then, a mask will be applied on the weighted sum of these selected prompts to create explicit pathways for cross-modal prompts recovery. The masked prompts then first process through intra-modal recovery to establish basic modality-specific patterns while introducing light cross-modal influence, followed by cross-modal that further integrates information across modalities through attention mechanisms and selective enhancement. The result is a set of recovered prompts that maintain balanced representations from both modalities.
@@ 4.1 Set Up
Dataset. Following Zhang et al. (2023), we conduct experiments on two widely-used datasets: VQA v2 (Goyal et al., 2017) and NExT-QA (Xiao et al., 2021). VQA v2 contains 1.1 million pairs of real-world images and human-written questions, while NExT-QA includes 52K annotated video-based question-answer pairs. For VQA v2, we structure the continual learning experiments into 8 sequential tasks for DI and CI, while 10 sequential tasks for QI. In QI, each task introduces 1 new question type. For CI, each task contains 10 unique object classes. Under DI, each task contains 5 subtasks and each subtask focuses on 2 different question types about the same 10 object classes. For NExT-QA, we structure the experiments into 7 sequential tasks. For QI, each task introduces 1 new question type. For CI, 4 tasks contain 11 object classes each, and 3 tasks contain 12 object classes each. For DI, each task contains unique question types with 5 subtasks, where each subtask addresses the same 16 object classes.
@@
Implementation Details. We construct MM-Prompt as illustrated in Fig. 2. The pretrained transformer backbone consists of 12 stacked blocks for both encoder and decoder, with each attention layer containing 12 attention heads. All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory. Following DualPrompt (Wang et al., 2022a), we further divide our question and visual prompts into General (G) and Expert (E) types that target different transformer layers. We also include a weight decay (Krogh and Hertz, 1991) on the $\mathcal{L}_{\text{inter}}$ to prevent degeneration of this alignment mechanism.
@@ 4.2 Experimental Results
In this section, we investigate and evaluate MM-Prompt against 9 outstanding and state-of-the-art methods, including 6 general CL methods Wang et al. (2022a); Smith et al. (2023); Wang et al. (2022b); Khattak et al. (2023); Menabue et al. (2024) and 3 CVQA methods Qian et al. (2023); Zhang et al. (2023); Cai and Rostami (2024). For fair comparison, all models contain at least two types of prompts(Q and V) except L2P* and use approximately the same number of prompts with identical injection layers if possible.
Based on Table 1, we draw several important conclusions: (1) Methods adopting cross-modal prompt isolation like Dual Prompt (Wang et al., 2022a) and L2P (Wang et al., 2022b) consistently show higher forgetting rates and lower accuracy across settings, confirming that isolated prompt selection and injection lead to degradation in performance due to imbalanced representations. (2) MM-Prompt consistently outperforms all the other methods across all settings. On VQA v2, it achieves 36.223% in DI (vs. MaPLe’s 35.187%), 39.240% in CI (vs. MaPLe’s 37.450%), and 16.757% in QI (vs. VQACL’s 15.525%). Notably, it reduces forgetting to less than half of the next best method in DI. On NExT-QA, MM-Prompt similarly outperforms in accuracy and maintains the lowest forgetting rates. While the second-best method varies across different incremental scenarios, MM-Prompt’s consistent superiority highlights its robustness across diverse settings. (3) We observe striking dataset-dependent difficulty patterns: VQA v2 (static images) shows poorest performance on QI, while NExT-QA (videos) struggles most with CI, suggesting that temporal dynamics fundamentally alter how modality conflicts during continual learning. (4) Among comparison methods, MaPLe (Khattak et al., 2023) performs relatively well, which we attribute to its gradient-based visual-linguistic interactions. (5) Specialized CVQA methods like Triplet (Qian et al., 2023) and VQACL (Zhang et al., 2023) underperform despite their success in classification-based settings and with memory buffers, highlighting MM-Prompt’s efficiency for maintaining balanced representations across modalities. These results collectively validate our approach’s effectiveness in maintaining balanced cross-modal representations across diverse incremental learning scenarios.
@@ 4.3 Modality Integration Analysis
To demonstrate that isolated prompts amplify modality imbalance, we train the models as normal and evaluate them on joint, vision-only and question-only inputs, then compute Modality Merge Effectiveness and Modality Difference, details for these settings are in Appendix A. As shown in Fig. 4, existing prompt-based methods (Wang et al., 2022a, b; Zhang et al., 2023) that adopt cross modal prompts isolation show low merge effectiveness and high modality gaps during the continual learning process. In contrast, our method maintains consistently high merge effectiveness and minimal modality disparity across new tasks, while achieving superior overall accuracy. These improvements stem from our two-component design: cross-modal prompt query prevents selection bias by enriching queries with complementary information, while cross-modal prompt recovery creates explicit pathways for modality interaction and fulfill them with cross-modal information. Together, these mechanisms counteract modality imbalance at critical points where isolation typically occurs in existing methods.
@@
Figure 5: Bottom up attention visualization(Anderson et al., 2018) during inference on Dual Prompt (Wang et al., 2022a) and $M$-Prompt.
(a) Dual PromptWang et al. (2022a) ($A:\text{33.601},F_{\text{inter}}:2.660$)
(b) $M$-Prompt ($A:\text{36.223},F_{\text{inter}}:\text{0.447}$)
@@
Figure 6: Visualization of injected prompts and input feature using t-SNE (Van der Maaten and Hinton, 2008).
@@
Figs. 5 and 6 provide further qualitative evidence. We visualize the attention from the model on the bounding boxes during the inference using (Anderson et al., 2018). The attention maps (Fig. 5) show MM-Prompt precisely focuses on relevant image regions (correctly identifying “broccoli” and “airport”), while existing methods such as Dual Prompt produce incorrect answers with diffuse attention patterns, indicating possible occurrence of modality bias (Ramakrishnan et al., 2018). The t-SNE visualization (Van der Maaten and Hinton, 2008) of injected prompts (Fig. 6) confirms the cross-modal merging effectiveness of MM-Prompt. As shown in Fig. 6(b), MM-Prompt’s prompts form more tightly assembled clusters compared to Dual Prompt in Fig. 6(a), where prompts remain clearly separated by modality. This observation demonstrates how our cross-modal prompt query and recover mechanisms help prompts encode integrated information across modalities rather than remaining isolated within modality-specific subspaces. Additionally, examining the input features from the question “What green vegetable is on the plate?” in Fig. 5 reveals another key difference: in MM-Prompt, input features position at the intersection between Q and V prompt clusters, while prompts in Dual Prompt remain distinctly separated by modality. This demonstrates that MM-Prompt’s injected prompts contains shared representational properties that facilitate effective cross-modal understanding and prevent modality-isolated prompts throughout sequential learning.
Methods NExT-QA
Mem AA ($\uparrow$) $F_{\text{inter}}$ ($\downarrow$)
Triplet (Qian et al., 2023) 500 27.996 3.282
Dual Prompt (Wang et al., 2022a) 500 27.641 3.765
CODA (Smith et al., 2023) 500 30.632 2.471
Maple(Khattak et al., 2023) 500 30.049 3.112
VQACL (Zhang et al., 2023) 500 30.753 2.692
CluMo (Cai and Rostami, 2024) 500 29.776 3.248
$M$-Prompt 500 32.394 2.259
$M$-Prompt 200 29.945 3.112
$M$-Prompt 1000 34.268 2.427
Figure 7: Time spent processing 100 samples.
(a) Dual Prompt (Wang et al., 2022a)
(b) $M$-Prompt
@@ 4.4 Ablation Study
Impact of Task Order. Figs. 8(a) and 8(b) demonstrate MM-Prompt’s stability across different learning scenarios. MM-Prompt maintains more consistent performance across different task orders, while Dual Prompt(Wang et al., 2022a) exhibits more varied performance. By maintaining balanced representations through explicit cross-modal pathways in injected prompts, MM-Prompt reduces the cascading effects of task order biases that typically compound in existing prompt methods. This order-invariant behavior strongly evidences that MM-Prompt successfully maintains balanced representations throughout the continual learning process.
@@
CQ CR AA ($\uparrow$) $F_{\text{inter}}(\downarrow)$
@@
Performance with Memory Buffer For NExT-QA dataset, which contains complex video question-answering tasks with longer sequences and fewer training data, we also evaluate models with a memory buffer of 500 samples (Table 2). As a result, MM-Prompt substantially outperforms all the other methods, given that all methods benefit from the same memory capacity. These results confirm that MM-Prompt’s approach to create explicit pathways for modality interaction remain effective when supplemented with traditional continual learning techniques (Rolnick et al., 2019; Buzzega et al., 2020).
Computational Efficiency. Fig. 7 shows that the inference efficiency of MM-Prompt remains competitive with existing methods. Processing 100 samples takes 0.179 seconds, which is more efficient than MAPLE (Khattak et al., 2023) (0.211s) and TRIPLET (Qian et al., 2023) (0.226s), while only slightly slower than Dual Prompt (Wang et al., 2022a) (0.158s). This modest computational overhead delivers substantial performance improvements and significant forgetting reduction compared to the Dual Prompt (Wang et al., 2022a). This demonstrates that fostering the cross-modal understanding can be achieved with reasonable computational costs, making MM-Prompt practical for real-world continual learning applications.
@@ A.1 Dataset Partitioning
For VQA v2 Goyal et al. [2017], we organize the continual learning experiments into 8 sequential tasks for both Class Increment (CI) and Dual Increment (DI) settings. In the CI setting, each task introduces 10 unique object classes. In the DI setting, each task consists of 5 subtasks, each focusing on 2 distinct question types about the same 10 object classes. For the Question Increment (QI) setting, we define 10 sequential tasks, each introducing 1 new question type.
For NExT-QA Xiao et al. [2021], we structure the experiments into 7 sequential tasks due to the dataset’s scale. In the QI setting, each task introduces 1 new question type. In the CI setting, 4 tasks contain 11 object classes each, while the remaining 3 tasks contain 12 object classes each. In the DI setting, each task consists of 5 subtasks with unique question types, each subtask cover 16 different object classes.
@@ A.2 Partitioning Details and Group Definitions
For NExT-QA [Xiao et al., 2021], we structure the experiments into 7 sequential tasks following the approach of Zhang et al. [2023]. In the DI setting, we implement the hierarchical structure defined in Zhang et al. [2023], with question increments ordered according to Table 8. For the QI setting, we utilize the question type ordering specified in Table 8 while maintaining consistent visual classes across all tasks. In the CI setting, we employ the object class groups outlined in Table 5 while preserving all question types throughout the tasks.
@@ B.1 Modality Merge Effectiveness and Modality Difference
$\text{Modality Merge Effectiveness}=\text{Acc}^{\text{joint}}-\max(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}}).$ (12)
@@
$\text{Modality Difference}=\max(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}})-\min(\text{Acc}^{\text{V-Only}},\text{Acc}^{\text{Q-Only}}).$ (13)
@@
Question Types
Recognition
Location
Judge
Commonsense
Count
Action
Color
Type
Subcategory
Causal
@@ B.2 Insertion Layers and Number of Prompts
Number of Prompts AA ($\uparrow$) $F_{\text{inter}}$($\downarrow$)
@@
Following Wang et al. [2022a], we insert General prompts at lower layers {1,2} and Expert prompts at intermediate layers {3,4,5}. This configuration follows the hierarchical processing pattern of the transformer, where lower layers handle more specific, fine-grained patterns, while intermediate layers process more abstract, transferable representations.